# Automating Sensor Characterization with Bayesian Optimization

### v1

@@ 3 Characterization and Optimization Methods
We optimize five operating parameters — four voltages and the number of CDS integration samples (see Table 1). Given the tightly coupled nature of the bias voltages and CDS, we expect these five parameters to have the greatest impact on the response of the detector.
We use BO to efficiently sample detector operational parameters (see Table 1) and identify optimal sets of operating parameters to maximize the SNR. At each step, the detector’s operating parameters are set. Then, the sensor is exposed to light, read out, and the objective value $F(x_{i})$ is calculated from the image.
Figure 3: Illustration of Bayesian optimization showing the true function (dashed), Gaussian process posterior mean and uncertainty (black line, shaded), observations (points), and the acquisition function (green) with its maximum (red triangle). Illustration is inspired by Fig. 1 in Ref. (B. Shahriari et al., 2016).
Figure 4: Left: An example image taken with this technique. The shading distinguishes regions of the image: prescan, active, overscan, and reverse overscan regions. Each region is used to diagnose an aspect of the detector’s performance. Right: Histograms of the pixel-charge distribution corresponding to the active (black) and reverse overscan (blue) regions highlighted by open rectangles in the left panel. The active pixel levels are distinct from those in the reverse overscan region, and this separation defines the signal $S$. The width of the reverse overscan represents the noise in the readout system NN.
The optimization routine generates an initial random sample of points in the five-dimensional parameter space. Then, during the BO stage, parameters are selected by maximizing an acquisition function, which encodes the tradeoff between exploring new regions of the parameter space and exploiting regions already known to perform well. At each iteration, the acquisition function’s maximal point defines the location of the next sampled parameters, and the observation (evaluated $F(x_{i}$)) reduces the uncertainty in that region. For each observation, the acquisition function and the model’s interpretation of the objective function’s landscape is updated. This process continues for the desired number of steps, $n$. Figure 3 shows an illustrative example of this process in a one-dimensional space.
Some parameter combinations result in detector artifacts, and the measured SNR does not correlate with charge sensitivity. Therefore, we incorporate the functions $P_{k}$ to penalize and suppress such effects. Therefore, the BO minimizes
$F(x_{i})=\frac{1}{\mathrm{SNR}(x_{i})}+\sum_{k=1}^{5}\alpha_{k}P_{k}^{\beta_{k}},$ (1)
where $x_{i}$ are the voltages and the CDS parameter, $P_{k}$ are polynomials designed to produce large values for images with saturation, non-uniformity in the active region, and artifacts from charge transfer effects, and $\alpha_{k}$ and $\beta_{k}$ are constants designed to scale the impact of $P_{k}$ on the objective function. A description of the the penalty functions are listed in Table 2.
$\alpha_{k}$ $\beta_{k}$ $P_{k}$ Description
$\frac{1}{2}$ $\frac{1}{3}$ $\max\!\left(0,\;-S\right)$ Negative Signal
$\frac{1}{2}$ $\frac{1}{3}$ $\max\!\left(0,\;N-S\right)$ Noise Exceeds Signal
$\frac{1}{3}$ $\frac{1}{3}$ $\max\!\left(0,\;2\,N-\sigma_{active}\right)$ Saturation
$\frac{1}{250}$ 22 $\max\!\left(0,\;U-1\right)$ Image Uniformity
11 11 $\phi(0;\sigma)-\phi(\mu_{\mathrm{reverse}};\sigma)$ Charge in Reverse Overscan
Table 2: Summary of the penalty used in the BO in Eqn. 1. The penalties are designed to suppress nonphysical or undesirable image properties: negative signal values, cases where noise dominates signal, and conditions that approach image saturation. They also enforce spatial uniformity by penalizing structure in the image that exceeds what is expected from random noise, quantified by the metric $U$ ($U\approx 1$ when the image is uniform). Excess charge in the reverse-overscan region is penalized using the Gaussian density function $\phi$, which compares the measured value to the distribution peak.
4 Experimental Results and Discussion
For the experimental setup for the BO application for sensor calibration, we first discuss the structure of the image acquired by the sensor and define signal and noise. The left panel of Figure 4 shows an image acquired during the experiment (during optimization), which reflects the standard CCD layout: a prescan, active region, overscan, and reverse overscan. The prescan provides a fixed electronic baseline at the start of the readout. The active region records the stimulus-dependent signal. The overscan (at the end of the readout) serves as a baseline diagnostic. The reverse overscan provides an additional reference that doesn’t have charge-transfer artifacts. The corresponding pixel-value distributions (right panel of Figure 4) highlight the distinct separation between signal and noise. The values in the active region are cleanly offset from those in the overscan, and the width of the reverse overscan distribution reflects the noise in the readout system. Since the CCD is uniformly illuminated, such separation is indicative of an optimal image and thus provides a useful diagnostic of both sensor performance and calibration quality.
Figure 5: Corner plot of the objective function landscape. Diagonal panels: one-dimensional partial dependences of the objective with respect to each parameter (others marginalized). Off-diagonal panels: pairwise partial dependences. The black points mark the configurations explored during burn-in and BO; dark purple colors represent higher objective values and bright yellow colors represent lower values (lower objective is better under the defined metric). The red vertical markers on the diagonal panels denote the coordinate-wise minimizers (argmin) of the partial dependence curves $PD_{j}(\theta_{j})$ over the explored range.
@@ 4 Experimental Results and Discussion
We used 500 iterations—150 randomly sampled points for initialization and burn-in, followed by 350 points guided by BO. Figure 5 displays the five-dimensional objective landscape. The diagonal panels report one-dimensional partial dependencies—the objective versus a single parameter $\theta_{j}$, marginalizing the others,
$PD_{j}(\theta_{j})=\mathbb{E}_{-\theta_{j}}[F(\theta_{j},\theta_{-j})],$
while the off-diagonal heatmaps show pairwise partial dependencies $PD_{jk}(\theta_{j},\theta_{k})$. Across these summaries, the optimizer consistently identifies $V_{DS},V_{GS},$ and $V_{Offset}$ as the dominant parameters, while CDS integration samples and the isolation-guard voltage $V_{I}$ exhibit weaker effects.
The structure of the landscape explains these patterns. In panels containing $V_{DS}$ (the first column of Figure 5), there is a pronounced gradient along the $V_{DS}$ axis across multiple pairings, indicating strong sensitivity to $V_{DS}$. Panels containing $V_{GS}$ (the second column) show vertical banding, indicating sensitivity to $V_{GS}$ largely independent of the paired parameter. The $V_{DS}-V_{GS}$ interaction does not present a clean oblique valley in this dataset; instead, it appears as banding, suggesting that both biases matter strongly but interact only weakly in the parameter ranges in the figure. The $V_{Offset}$ diagonal shows a broad, shallow ‘U’ shape, consistent with placing the baseline comfortably within the ADC window while avoiding penalty regions. By contrast, the CDS on the diagonal of the figure is monotonically decreasing across the tested range (more integration samples reduce the contribution of read noise without revealing an interior optimum) and the isolation-guard $V_{I}$ diagonal exhibits mild curvature. To make these weaker trends visible, the vertical scales of the CDS and $V_{I}$ marginals are expanded. The black samples cluster in lower-objective bands of these maps (as expected under BO), which should be read as concentration near favorable regions. These patterns align with the image-level definitions used earlier: regions of lower objective (per the colorbar) correspond to settings where the active region distribution is well separated from the reverse-overscan (large signal) and the reverse-overscan width (noise) is small, with penalty terms steering away from saturation, non-uniformity, or transfer-artifact regimes. Finally, the relative importance of the parameters remains sensor- and objective-dependent: for SiSeRO-class devices, the volatility of the output signal and the front-end electronics elevates $V_{DS},V_{GS},$ and $V_{Offset}$; under different goals, CDS and $V_{I}$ can become decisive.
@@
Finally, Figure 6 shows the running minimum of the objective $F$ after $n$ calls during the 500-step experiment (150 random burn-in, 350 BO-guided). The curve shows rapid improvement early on, followed by steadier gains as the optimizer refines near-favorable regions.
@@ 5 Conclusions and Outlook
Looking forward, the proposed technique has already been successfully validated on a Skipper CCD, confirming its robustness across different sensor architectures. We are extending its use to additional novel sensors (A. J. Lapi et al., 2024) and will establish it as a standard procedure in our laboratory, streamlining the characterization of new devices as they are fabricated and packaged. Ultimately, our objective is to generalize from single-sensor optimization to the automated calibration of full sensor arrays, providing a scalable framework for next-generation detector systems.

### v2

@@ 4 Optimization Experiments: Results and Discussion
We used 500 iterations—150 randomly sampled points for initialization and burn-in, followed by 350 points guided by BO. For each iteration after burn-in, the objective function is calculated and incorporated into the surrogate model, which guides the choice of the next point. Figure 5 displays the objective function in single dimensions (diagonal panels) and in two dimensions (off-diagonal panels) across the five-dimensional parameter space. One-dimensional partial dependencies are the objective $F(x_{i})$ versus a single parameter $\theta_{k}$, marginalizing the others:
$PD_{k}(\theta_{k})=\mathbb{E}_{-\theta_{k}}[F(\theta_{k},\theta_{-k})].$
Across the one-dimensional partial dependencies, the BO procedure consistently identifies strong dependence on $V_{DS},V_{GS},$ and $V_{Offset}$, while CDS integration samples and the isolation-guard voltage $V_{I}$ exhibit weaker effects. The off-diagonal heatmaps show pairwise partial dependencies $PD_{kl}(\theta_{k},\theta_{l})$, with the black points showing the parameter values explored during burn-in and BO. In the pairwise partial dependence maps, the color scale represents the objective function marginalized over the remaining parameters, with higher objective values shown in dark purple and lower values in bright yellow. We note that the apparent overlap of sampled configurations with higher-objective regions in some two-dimensional projections does not imply that those configurations themselves yielded high objective values; rather, it reflects the projection of a five-dimensional optimization onto a two-dimensional subspace.
The panels in the first column of Figure 5 show the sensitivity of the objective function to $V_{DS}$ alone and when correlated with other parameters. The objective function exhibits a broad minimum and maximum in $V_{DS}$, indicating apparent but low sensitivity to that parameter. In the $V_{DS}-V_{GS}$ correlation, there are several small regions of maxima in the objective function, indicating relatively high sensitivity to those parameters in tandem. This is expected because $V_{GS}$ and $V_{DS}$ set the operating point of the SiSeRO transistor, affecting the sensitivity and noise of the amplifier. In the $V_{DS}-V_{Offset}$ correlation, there are two relatively large regions of local maxima, indicating relatively low sensitivity to those parameters in tandem. There is a nearly monotonic gradient along the $V_{DS}$ axis with respect to CDS and $V_{I}$.
Next, the second column shows the degree and type of sensitivity to $V_{GS}$ alone and when correlated with three other parameters. In the $V_{GS}-V_{offset}$ correlation, partial vertical bands exist. In the correlations with CDS and $V_{I}$, there are full vertical bands. In the third column, the objective function with respect to $V_{Offset}$ (diagonal) shows a broad, shallow ‘U’ shape; this is consistent with placing the baseline signal level comfortably within the ADC window while avoiding penalty regions. In the fourth column, the CDS (diagonal) monotonically decreases because increasing the number of integration samples reduces read noise without introducing any adverse detector artifacts. The objective function exhibits mild curvature with respect to the isolation-guard voltage $V_{I}$.
Taken together, the partial dependence plots indicate that parameters exhibiting larger variation in their one-dimensional partial dependence, most notably the MOSFET bias voltages ($V_{GS}$ and $V_{DS}$) and the offset voltage ($V_{Offset}$), play a dominant role in shaping the objective landscape under the present metric. The distributions of sampled configurations reflect this behavior, with non-uniform sampling patterns in the $V_{GS}$, $V_{DS}$, and $V_{Offset}$ pairwise dependences, while parameters with comparatively flat one-dimensional partial dependence show no visually distinct preference. This hierarchy of parameter influence is expected to depend on both the sensor architecture and the optimization objective: for SiSeRO-class devices, the volatility of the output signal and coupling to the front-end electronics emphasize transistor biasing and ADC conversion parameters, whereas other detector architectures or optimization goals may place greater weight on CDS settings or $V_{I}$.
@@
Figure 6 shows the minimum of the objective value $F$ after $n$ optimization steps; this includes both the burn-in and the BO-guided phases. The curve shows rapid improvement early on, followed by steadier gains as the optimizer refines near-favorable regions.
@@ 5 Conclusions and Outlook
Looking forward, the proposed technique has already been successfully validated on a Skipper CCD, confirming its robustness across different sensor architectures. We are extending its use to additional novel sensors [9] and will establish it as a standard procedure in our laboratory to streamline the characterization of new devices as they are fabricated and packaged. Ultimately, our objective is to generalize from single-sensor optimization to the automated calibration of full sensor arrays, providing a scalable framework for next-generation detector systems.
Acknowledgments